OpenStudy (anonymous):

Use the Quadratic formula to solve the equation. Simplify any radicals...x2=7-4x

OpenStudy (anonymous):

if you bring everything to one side to make it simpler it would be x^2 + 4x - 7 =0

OpenStudy (anonymous):

now just plug that into the quadratic formula. do you have the formula memorized?

OpenStudy (anonymous):

i do

OpenStudy (anonymous):

alright, so what would the quadratic formula for that equation look like?

OpenStudy (anonymous):

how do i right it in?

OpenStudy (anonymous):

-b+-sqare root b2-4ac all over 2a

OpenStudy (anonymous):

first lets find out what our a, b, and c are: since we have x^2 by itself we know a = 1 since we have 4x as our second term our b = 4 and -7 is our last term so we have c = -7 so now plug that into your formula

OpenStudy (anonymous):

-2+-sqare root 11?

OpenStudy (anonymous):

you should get something that looks like this: \[-4\pm \sqrt{(4)^2 - 4(1)(-7)} \over 2(1)\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

yeah, you got it! -2 +- sqrt(11) is right! i'm pretty impressed, most people who come here just want answers, you have my respect good sir/madam

OpenStudy (anonymous):

thanks! I am trying to get them right! =)

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (anonymous):

well you are doing a good job of it so far!! :)

OpenStudy (anonymous):

no problem! glad i could be of service!

OpenStudy (anonymous):

i just thought everything had to be positive thats where i was getting confused

OpenStudy (anonymous):

nope, you are allowed to have negatives wherever you want, it just starts getting funky when they are underneath the square root. but i'm sure you'll learn about that soon enough. :)

OpenStudy (anonymous):

if i have another question do i ask it here or do i repost it?

OpenStudy (anonymous):

you can ask here if you like, or repost, it's really up to you. i'd be glad to help you with another one either way.

OpenStudy (anonymous):

ok cool thanks!...its a complex fraction...

OpenStudy (anonymous):

bring it on! lol

OpenStudy (anonymous):

\[9 + 3/x \over x/4 + 1/12\]

OpenStudy (anonymous):

so far i have all the work down to \[12(9x+3) \over x(3x+1) \] and then I get lost...

OpenStudy (anonymous):

are you just supposed to be reducing it?

OpenStudy (anonymous):

the directions are to simplify..

OpenStudy (anonymous):

ok, sounds good, you are almost done actually! the trick here is to get rid of the (9x+3) and (3x+1)

OpenStudy (anonymous):

i have 9x+ 3 over x Multiplied by 12 over 3x+1?

OpenStudy (anonymous):

you can do this by factoring a 3 out of the (9x+3) which will give you 3(3x+1) since you have the same thing on the bottom you can cancel both of them

OpenStudy (anonymous):

so it it 36 over x?

OpenStudy (anonymous):

so it will look like: \[12(3)(3x+1) \over x(3x+1)\]

OpenStudy (anonymous):

yup, when you cancel and multiply that is exactly what you get.

OpenStudy (anonymous):

ok I was ending up with an extra 3 thats why I wasn't sure! thanks

OpenStudy (anonymous):

no problem! glad i could help you work that one out.

OpenStudy (anonymous):

I have a couple more if you're not to busy...